If 35 000 accumulates to 48 328 at a continuous compounded rate of year then the term under consideration is


1
Expert's answer
2022-03-08T13:33:02-0500

Solution


Using the Formula


"A=Pe^{rt}"


Here "A" is the amount accumulated, "P" is the principal amount invested, "r" is the interest rate and "t" is the tenure.


Therefore, we can write,


"48328=35000(e^{rt})"


"e^{rt}=\frac{48328}{35000}"


"e^{rt}=1.3808"


"ln(e^{rt})=ln(1.3808)"


"rt\times ln(e)=ln(1.3808)"


"rt\times (1)=ln(1.3808)"


"t=\frac{ln(1.3808)}{r}"


Therefore, knowing the value of "r", we can find the tenure.


For example if "r=6.8\%"


Then


"t=\frac{ln(1.3808)}{6.8\%}"


"t=\frac{ln(1.3808)}{0.068}"


"t=4.745"  years